The dielectric of a parallel plate capacitor is to
Answer #1

K =4.6 , A =0.5 m^2, C =0.1 uF

(A) C = AKeo/d

0.1*10^-6 = (0.5*4.6*8.85*10^-12)/d

d=0.204 mm

(b) V =12 V

Q = VC = 12*0.1 = 1.2 uC